OpenStudy (anonymous):

integrate sqrt((x^3 -3)/x^11)

OpenStudy (blurbendy):

does that square root apply to the whole function

OpenStudy (blurbendy):

or just the top part

OpenStudy (blurbendy):

fraction*

OpenStudy (anonymous):

whole

hartnn (hartnn):

is this your integral ? \(\int \sqrt{\dfrac{x^3-3}{x^{11}}}dx\)

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

do you have an answer? because i think i have one

OpenStudy (anonymous):

yes \[(2/27)\sqrt{(1-x^3)^3}\]

OpenStudy (anonymous):

ops wrong

OpenStudy (anonymous):

\[(2/27)\sqrt{(1-3/x^3)^3}\]

OpenStudy (blurbendy):

close, after (2/27) i have x^12 and then i have ( (x^3 - 3) / (x^11) )^3/2

OpenStudy (anonymous):

they are the same :) but how ?

OpenStudy (abb0t):

fancy squiggly line.